OpenStudy (anonymous):

Find the unit tangent, principal normal,and wrtie an equation in x,y,z for the osculating plane at the point on the curve that corresponds to the indicated value of t. 35. r(t) = i + 2tj +t^2k ; t =1

OpenStudy (amistre64):

the normal to a curve is the gradient of the function

OpenStudy (amistre64):

which sounds odd when the brits call gradient a slope ...

OpenStudy (amistre64):

but we have a vector function, not a surface right

OpenStudy (anonymous):

yes r(t) is a vector

OpenStudy (amistre64):

we still need r', and then unit it

OpenStudy (amistre64):

then the derivative of the unit tangent produces the normal ....

OpenStudy (amistre64):

crossing the tangent and normal gives us the ... I know they call it a B vector, curve maybe?

OpenStudy (anonymous):

so the unit tangent is r'(t)/ its magnitude?

OpenStudy (amistre64):

yes r = 1 i + 2t j +t^2 k r' = 0 i + 2 j +2t k |r'| = sqrt(4+4t^2) = 2sqrt(1+t^2)

OpenStudy (amistre64):

for some reason, the normal has to be calculated using the unit tangent ...

OpenStudy (anonymous):

ok, so we include the t when calculating the magnitude

OpenStudy (amistre64):

of course

OpenStudy (anonymous):

|r'| = sqrt(4+4t^2) = 2sqrt(1+t^2) how did you take out the 2?

OpenStudy (amistre64):

\[T = 0~ i + \frac{2}{2\sqrt{1+t^2~}}~ j +\frac{2t}{2\sqrt{1+t^2}}~k\] well, factor out the 4, and by now im well aware that sqrt4 = 2

OpenStudy (anonymous):

oh i see now.

OpenStudy (amistre64):

sqrt(4+4t^2) sqrt[4(1+t^2)] sqrt(4) sqrt(1+t^2) 2sqrt((1+t^2)

OpenStudy (anonymous):

ok. so then we take the derivitive of this to find the normal? or am i misunderstanding

OpenStudy (amistre64):

N = T'/|T'|

OpenStudy (amistre64):

taking the derivative, that right

OpenStudy (anonymous):

so how would you go from T(t) = 1/2sqrt(1+t^2)(2j +2tk) to its derivitive

OpenStudy (amistre64):

\[T = 0~ i + \frac{2}{2\sqrt{1+t^2~}}~ j +\frac{2t}{2\sqrt{1+t^2}}~k\] \[T' = 0~ i + \left(\frac{1}{\sqrt{1+t^2~}}\right)'~ j +\frac12ln|\sqrt{1+t^2}~|~k\] that j part looks maybe like a tangent?

OpenStudy (amistre64):

nah, ln(sec+tan) i think

OpenStudy (amistre64):

lol, im integrating

OpenStudy (anonymous):

yah you had me confused for a second.

OpenStudy (anonymous):

so basically need the derivitive of 1/sqrt(1+t^2) and derivitive of t/sqrt(1+t^2)

OpenStudy (amistre64):

taking the derivative .... \[T' = 0~ i - \frac{4t}{(1+t^2)^{3/2}~}~ j +(...)~k\] yeah

OpenStudy (amistre64):

ive got to run, but im sure you can manage from here :) take the normal, and anchor it ti r(t), at t=1 to define the plane

OpenStudy (anonymous):

ok thank you